OpenStudy (anonymous):

ln0.25x/ln1.5 = 0.75 solve for x

OpenStudy (ranga):

Multiply both sides by ln1.5 Then use the formula If ln(y) = k then y = e^k

OpenStudy (anonymous):

how do I do that?

OpenStudy (ranga):

ln0.25x/ln1.5 = 0.75 Multiply both sides by ln1.5 ln0.25x = (0.75)(ln1.5) = 0.3041 0.25x = e^(0.3041) = 1.3554 Divide both sides by 0.25 x = 5.4216 To double check redo the calculations with your calculator.

OpenStudy (anonymous):

ty its correct!

OpenStudy (ranga):

Cool. You are welcome.